Write an equation of the line that is parallel to y=5x+4 and passes through the point -4,3.
Answer by TimothyLamb(4379)

Parallel lines have equal slope.
---
Given the line L1 in slope-intercept form: y = 5x + 4
its slope is: 5
---
To find a line L2 parallel L1, and passing through the point p2 = (-4, 3),
use the point-slope form: y - y1 = m(x - x1)
---
y - 3 = 5(x - -4)
y - 3 = 5x + 20
y = 5x + 20 + 3
---
Answer, an equation for line L2 is:
y = 5x + 23
---
Check the work by plugging-in point p2 into the equation:
---
3 = 5(-4) + 23
3 = -20 + 23
3 = 3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php